omg123456:

If the height of the cylinder is doubled and the diameter is tripled, by what factor will the volume of the cylinder change?

jhonyy9:

radius of cylinder let r and the height let h volume of cylinder equal area of base time height = pi*r^2 *h = V_1 with height doubled so 2h and diameter tripled what mean 2r tripled so 3*2r = 6r the volume in this case equal pi *(6r)^2 *2h = pi*36r^2 *2h = 72*pi*r^2 *h = 72*V_1 hope helped understandably

sillybilly123:

🥳 https://www.youtube.com/watch?v=sHKUcXNq_xk